OpenStudy (abbycross167):

At Generic High School, in the year 2001, there were 20 students per every computer. In 2008, there were 4.5 students per computer. Write two ordered pairs to represent the ratio of student to computer. USE X=0 FOR THE YEAR 2000. ~ Can someone please help me answer this?

OpenStudy (mathstudent55):

If x = 0 represents the year 2000, then what is x for year 2001?

OpenStudy (abbycross167):

1?

OpenStudy (mathstudent55):

Correct. In other words, x = year - 2000 For 2000, x = 2000 - 2000 = 0 For 2001, x = 2001 - 2000 = 1 For 2002, x = 2002 - 2000 = 2 Etc. What is x for year 2008?

OpenStudy (abbycross167):

8?

OpenStudy (mathstudent55):

Correct. So far we have the two x-coordinates of the two ordered pairs we need: \((1, ~~~)\) and \((8, ~~~)\) Now we need the corresponding y-coordinates. When x = 1, in the year 2001, what was the ratio of students to computers? In 2008, when x = 8, what was the ratio of students to computers? Those are the y-coordinates you need to fill in those ordered pairs above.

OpenStudy (abbycross167):

OK @mathstudent55 umm I'm kind of confused (I'm not very good at math I'm sorry) What do you mean "what was the ratio of students to computers?" like this~ 4.5 students per computer?

OpenStudy (mathstudent55):

Yes. You are told the ratio of students to computers was 20 in 2001 and 4.5 in 2008.

OpenStudy (abbycross167):

oh ok, thank you very much!!

OpenStudy (mathstudent55):

You're welcome. Did you get \((1, 20)\) and \((8, 4.5) \) ?

OpenStudy (abbycross167):

yes sir